How many grams of NO are needed to produce 20g of NO2?
1
Expert's answer
2020-12-08T05:33:17-0500

Solution:

Nitric oxide (NO) reacts with oxygen gas (O2) to form nitrogen dioxide (NO2):

2NO(g) + O2(g) → 2NO2(g)

According to the equation: n(NO) = n(NO2)


Moles of NO2 = n(NO2) = Mass of NO2 / Molar mass of NO2

The molar mass of NO2 is 46 g/mol.

Hence,

n(NO2) = 20 g / 46 g mol-1 = 0.4348 mol


n(NO) = n(NO2) = 0.4348 mol

Moles of NO = n(NO) = Mass of NO / Molar mass of NO

The molar mass of NO is 30 g/mol.

Hence,

Mass of NO = Moles of NO × Molar mass of NO

Mass of NO = 0.4348 mol × 30 g/mol = 13.044 g

Mass of NO is 13.044 grams


Answer: 13.044 grams of NO are needed to produce 20g of NO2.


(20 g NO2) × (1 mol NO2 / 46 g NO2) × (30 g NO / 1 mol NO) = 13.043 g NO